# Circle packing on spherical caps

### v1

@@ Abstract
We have carried out an extensive search for the densest packing configurations of congruent disks on spherical caps of selected angular widths ($\theta_{\rm max}=\pi/6$, $\pi/4$, $\pi/2$, $43\pi/$ 4, $65\pi/$ 6) and for several values of $N$. The numerical results obtained in the present work have been used to establish (at least qualitatively) some general features for these configurations, in particular the behavior of the packing fraction as function of the number of disks and of the angular width of the cap, or the nature of the topological defects in these configurations (it found that as the curvature increases the overall topological on the border tends to become more negative).
@@ 1 Introduction
After all one should recognize that the eye of an insect is just a Nature–made, very efficient detector where the close packing of sensors on the curved surface has been the result of a long evolutive process and not of a computational algorithm. As a matter of fact there has been a consistent effort in recent years in developing ”bio-inspired” hemispherical cameras, drawing inspiration from the insect’s eyes [65, 66, 67, 68, 69, 70, 71, 72]. Finally, another problem where circle packing on a spherical cap is relevant is to the construction of geodesic domes in engineering and architecture [73, 74, 75]. A very relevant reference in this regard is the nice book by Popko [76].
@@ 2 The method
A major limitation of the original algorithm of ref. [80] and also of the algorithms discussed in refs. [81, 79, 82] is represented by the fact that it works only for special domains, such as the square, the circle or, more in general, regular polygons. This occurs because the algorithm does not work directly with the physical container (domain) of the disks, but with a smaller domain of the charges. Because of this the two domains must be related by a scale transformation, otherwise the nature of the problem would vary during the process and any comparison between configurations would not make any sense.

### v2

@@ Abstract
We have carried out an extensive search for the densest packing configurations of congruent disks on spherical caps of selected angular widths ($\theta_{\rm max}=\pi/6$, $\pi/4$, $\pi/2$, $43\pi/$ 4, $65\pi/$ 6) and for several values of $N$. The numerical results obtained in the present work have been used to establish (at least qualitatively) some general features for these configurations, in particular the behavior of the packing fraction as function of the number of disks and of the angular width of the cap, or the nature of the topological defects in these configurations ( it was found that as the curvature increases, the overall topological charge on the border tends to become more negative).
@@ 1 Introduction
A second exception is ref. [26], where solutions for $N=2,3,4,5,6,7,8$ have been conjectured on the basis of a computer analysis.
@@
One should recognize that an insect’s eye is a highly efficient, Nature-made detector, where the close packing of sensors on a curved surface is the result of a long evolutive process, not of a computational algorithm. As a matter of fact there has been a consistent effort in recent years in developing ”bio-inspired” hemispherical cameras, drawing inspiration from the insect’s eyes [66, 67, 68, 69, 70, 71, 72, 73]. Finally, another application of circle packing on a spherical cap is in the design of geodesic domes for engineering and architecture [74, 75, 76]. A key reference on this topic is the book by Popko [77].
@@ 2 The method
The primary limitation of the original algorithm presented in Ref.[81], as well as the algorithms discussed in Refs.[82, 80, 83], is that they are restricted to specific geometrical domains, such as squares, circles, or regular polygons. This limitation arises because the algorithm does not work directly with the physical container but with a smaller domain. Consequently, a scale transformation must be applied to relate the two domains, otherwise the nature of the problem would change during the process, rendering any comparison between configurations meaningless.
@@ 4 The fate of CHP
The analogous cases of $N=61$ and $N=91$ are plotted in the Figs. 16, 16, 18, 18. Similarly the packing density for $N=12$ and $N=31$ (non CHP configurations) are plotted in Figs. 19 and 20.